# Rethinking Positive Pairs in Contrastive Learning

### v1

@@ Abstract
We present Hydra, a universal contrastive learning framework for visual representations that extends conventional contrastive learning to accommodate arbitrary pairs. Our approach is validated using IN1K, where 1K diverse classes compose 500,500 pairs, most of them being distinct. Surprisingly, Hydra achieves superior performance in this challenging setting. Additional benefits include the prevention of dimensional collapse and the discovery of class relationships. Our work highlights the value of learning common features of arbitrary pairs and potentially broadens the applicability of contrastive learning techniques on the sample pairs with weak relationships.
@@ 1 Introduction
Figure 1: Similarity distribution of class pairs under the subspaces of feature extracted by SimCLR and Hydra for snake-lamp. We can hardly find any common visual features from the example of a garter snake and a table lamp, However, we find that the 500 dimensions with the lowest variance from SimCLR’s representation can separate snake-lamp from other classes. Hydra learns such a subspace while representation learning.
@@
We present Hydra by extending CL methods to accommodate arbitrary pairs, e.g., SimCLR. To show the robustness of dealing with disparate pairs, we train Hydra on ImageNet (Deng et al., 2009), with 1,000 diverse classes forming 500,500 class pairs, most of which are disparate. The positive pairs are randomly selected from the 500,500 class pairs. Our model achieves superior performance in the challenging situation, demonstrating its robustness and effectiveness in discovering common features. In summary, our contributions are as follows:
  • We propose Hydra a universal contrastive learning framework to allow learning visual representations from samples belonging to arbitrary class pairs through feature selection.
@@ 3.3 Universal Contrastive Learning
Figure 2: Universal contrastive learning for arbitrary class pairs. The feature filter generates a gate vector to activate the common features for the given class pair by averaging their label embeddings. Hydra learns visual representations by maximizing the agreement of common features between disparate samples (Hydra-Lamp) in the corresponding subspace.
@@
The feature filter is a vital component of Hydra to allow universal contrast learning. It controls the activation of each feature for a class pair y1-y2. The feature filter consists of two parts: label embedding and MLP layer. Label embedding converts discrete labels to continuous vectors. We use the mean of the two label vectors to represent the common information among a class pair. An MLP layer generates gate values to select the corresponding dimensions for the common information. The calculation is
@@ 4.1 Experimental Setting
All experiments are conducted using ResNet50 (He et al., 2016) as the backbone architecture, with ImageNet-1K (IN1K) as the training dataset. For baseline comparisons, we use pretrained models from official repositories: MoCov3, DINO, BarlowTwins, BYOL, and GPaCo; SimCLR implementation from VISSL (Goyal et al., 2021); Supcon trained for 600 epochs. Our proposed method is implemented in two variants: Hydra is built upon SimCLR framework, Hydra-MoCo is built upon SimCLR framework MoCov3. Detailed training configurations and hyperparameters are provided in appendix D.
@@ 4.2 Main Result
We adopt the transfer learning performance with linear prob to evaluate the quality of learned representations. Table 1 compares UCL to ICL and CCL in terms of their transfer learning performance on six classification tasks: Cars, CIFAR10, CIFAR100, FLW, Pets, and STL, with the average (AVG) score also provided. For a fair comparison, we obtain two groups of results by using momentum or not, and pick their representatives: SimCLR as ICL and Supcon as CCL without momentum, MoCov3 as ICL and GPaCo as CCL with momentum. Hydra demonstrate competitive performance in each group. Especially, Hydra MoCo as well as MoCoV3 stands out with the highest average scores of 94.0%, indicating their superior performance across all tasks. Notablely, Hydra MoCo is trained for 300 epochs, using only 30% of training for MoCov3. Overall, the table underscores the importance of uCL in achieving optimal transfer learning performance.
@@
Vision transformers (Dosovitskiy, 2020) are dominant in many CV tasks, especially beneficial from scaling in larger data (Dehghani et al., 2023). We study the effectiveness of our model for increasing the model size in table 2. We train 100 epochs for ViT-tiny, 300 epochs for ViT-small and ViT-base. Hydra outperforms both supervised and unsupervised learning methods for the ViT-tiny model. The experimental results demonstrate that Hydra scales well with increasing model size, outperforming or matching supervised learning across different ViT architectures. This scaling behavior, combined with Hydra’s ability to learn from arbitrary pairs, makes it a promising approach for leveraging large-scale, diverse datasets in visual representation learning tasks.
@@ 4.3 Universal Contrast Avoids Dimensional Collapse
To thoroughly investigate the dimensional collapse phenomenon, we construct IN1P, a focused subset of IN1K containing 12K images from 10 highly-related dog breeds. This smaller dataset allows us to extensively train on all possible class pairs, which is challenging in IN1K. We train models for 20,000 epochs to ensure comprehensive learning of all pairs. As shown in fig. 3(a), we compare the KNN classification performance (top-1 accuracy) on CIFAR10 across different methods. SimCLR exhibits clear signs of dimensional collapse after 3,000 epochs, evidenced by declining performance and decreasing singular values (Jing et al., 2022). This reveals that while the non-linear projector can delay dimensional collapse, it cannot prevent it entirely. In contrast, Hydra demonstrates sustained performance improvement throughout the extended training period, surpassing both Supcon and SimCLR. This robustness against dimensional collapse can be attributed to our framework’s ability to leverage rich common features between classes.
@@
Figure 3: Hydra prevents dimensional collapse and benefits from longer training. Trained on IN1P and evaluated on CIFAR10. Lower singular values suggest that the learned representations are concentrating information in fewer dimensions, that is, dimensional collapse.
@@ 4.4 Visualization
We analyze the quality of learned representations by examining embeddings from the backbone network (excluding the projector and filter). First, we compare the inter- and intra-class similarity distributions on the IN1K validation set. As shown in fig. 4, Hydra achieves the sharpest inter-class similarity distribution and the smallest overlap between inter- and intra-class distributions compared to Supcon and SimCLR. This indicates that Hydra learns representations that better preserve class-specific information while maintaining clear class boundaries.
To visualize these embedding properties, we apply $t$-SNE (Van der Maaten and Hinton, 2008) to the features of 10 classes from the IN1K validation set (fig. 5). The visualization reveals two key findings: (1) In the global space, Hydra maintains clear class separation as good as the methods optimized in the global space, even through our model is optimized in subspaces; (2) When examining the subspace for specific class pairs (e.g., Garter snake-Chihuahua), our model successfully identifies shared features that bring these seemingly disparate classes closer while preserving the overall structure of the embedding space. This demonstrates our model’s unique ability to learn both class-discriminative features and shared characteristics across arbitrary class pairs.
Figure 4: Similarity distribution for Hydra, Supcon, and SimCLR. The number in the middle denotes the overlap of two distributions. Smaller value means better class-separation.
@@
Figure 5: $t$-SNE visualization of learned embeddings for two contrasting groups: 5 dog breeds and 5 snake species. Large markers indicate class centers, with stars representing dog classes and points representing snake classes. While classes remain well-separated in the global space (abc), Hydra can selectively bring disparate classes closer in their designated subspace (d) through learned feature filtering.
@@ 5.1 Ablation
Figure 8: Impact of semantic distance between positive pairs (N-closest classes, N=0 means same class). We report KNN classification accuracy on CIFAR10. While Supcon’s performance drops significantly with increasing semantic distance, Hydra maintains robust performance through adaptively learning in subspaces.
@@
To systematically evaluate the limitations of supervised contrastive learning (Supcon) and identify scenarios where our method excels, we conduct experiments on IN25, a dataset of 25 classes with well-defined semantic relationships (details in section D.1). We systematically vary the semantic distance of positive pairs by composing them from N-closest classes, where N=0 represents pairs from the same class. As shown in fig. 8, we compare Supcon and Hydra on CIFAR10 transfer learning performance as we increase the semantic distance between positive pairs. Supcon’s performance degrades consistently as the semantic distance increases, highlighting its limitation in learning from disparate pairs. In contrast, Hydra maintains robust performance across different semantic distances, thanks to its ability to create appropriate subspaces for each pair.
@@ 5.2 Hyperparameter
We investigate the impact of each component in Hydra. To evaluate the quality of learned representations, we apply the widely used linear prob (Chen et al., 2020) on IN1K, KNN (K=10) on CIFAR10 as a reference of transfer learning. Hydra was trained on IN1K for 100 epochs. Table 3 demonstrates the effects of various hyperparameters. Our default settings for training are highlighted in Gray.
@@
We find that normalization at the end of the projector is critical to train Hydra. Without normalization, the training process becomes unstable, resulting in NaN errors. Conversely, the model barely converges with Batch Normalization (Ioffe and Szegedy, 2015). This is because Batch Normalization enforces each dimension to be informative (high variance), conflicting with our filter module, which selectively closes certain dimensions. Our model only functions effectively with Layer Normalization  (Ba et al., 2016).
@@
Similar to other contrastive learning methods (Chen et al., 2020; He et al., 2020), Hydra’s performance is influenced by the number of dimensions and the temperature parameter.
@@ 6 Conclusion
In this paper, we have reconsidered the concept of positive pairs in contrastive learning and explored the learning framework for arbitrary pairs. Our research breaks the limitation of positive pairs and demonstrates the potential of learning common features from seemly unrelated class pairs. The search space of positive pairs for different contrastive learning approaches. As the allowed distance for positive pairs increases, the search space expands exponentially. This expansion highlights the potential for more comprehensive and nuanced representation learning. Our approach, Hydra, demonstrates remarkable robustness in dealing with arbitrary pairs, even in a challenging case where positive pairs are randomly picked and most of them are disparate. The method successfully learns transferable representations under these challenging conditions, validating our core idea.
@@ Appendix A Understanding from A Case Study
Considering the large number and diversity of classes in IN1K, we use a case study to demonstrate how Hydra interprets the similarity between classes. Specifically, we select 17 classes belonging to the super synset ‘snake.n.01’, 16 classes belonging to the super synset ‘wading bird.n.01’, and 21 classes belonging to the super synset ‘furniture.n.01’ from the IN1K validation set. Each class contains 50 images. This case study demonstrates that Hydra promotes
We summarise the gate values of Hydra in Figure 9. Specifically, we pass 1K classes to the filter to generate 1000 vectors with 256-D gate values, which indicate the activation of the corresponding dimensions for each class. Each light point in the heatmap represents the activation status of a specific dimension for the corresponding class. From the figure, we can observe that the gates exhibit a roughly binary behavior (either activated or not), and the activation vectors for different classes are distinct. This indicates that each class creates a unique subspace within the feature space. These experimental observations support our hypothesis that Hydra can effectively learn to create class-specific subspaces.
@@
We investigate how Hydra identifies common features between similar classes within a coarse category. For the subclasses within a coarse class, the common features are dominant, i.e. showing high similarity in the global space, as discussed earlier. To delve deeper, we select the top-10 dimensions with the highest activation for subclasses within each coarse class, as illustrated in Figure 11. These dimensions likely represent the most salient features for distinguishing the subclasses. We then calculate the class similarity of features in the subspace defined by these dimensions for each coarse class. Figure 10 visualizes the class similarity in the subspaces computed for three coarse classes: Snake, Bird, and Furniture. Compared to the class similarity in the global space, the discriminative information is strengthened for the corresponding subspace. Specifically, the gap between intra and inter coarse-class similarity is magnified to 0.5 for Snake in the Snake-Snake Subspace, 0.7 for Bird in the Bird-Bird Subspace, and 0.5 for Furniture in the Furniture-Furniture Subspace. The features in the subspaces must contain distinct information about the class pair. These features also contain visual information about other classes such that these classes are still discriminative in the unrelated subspaces. These results demonstrate Hydra’s ability to create meaningful subspaces that amplify the differences between the classes, even when no common features appear to exist in the global space, like Furniture-Snake.
Figure 10: Class similarity in subspaces. The gap between intra and inter coarse-class similarity is widened in their corresponding subspaces, highlighting Hydra’s effectiveness in creating discriminative subspaces.
@@
We investigate three similarity measurements to evaluate the ability to identify the relationships between classes: 1) Dot product of gates: We calculate the gates for each class and sharpen the values to 1 if above 0.5 and to 0 otherwise. This measure indicates how similar the activated dimensions are between classes, reflecting the structural similarity of their subspaces. 2) Features in the global space: For each class pair, we use the global features to calculate the cosine similarity between all sample pairs coming from each class pair. We then use the average to denote the similarity between classes. This represents the overall visual similarity in the original feature space. 3) Class similarity in label embedding: We obtain the class vectors through the label embedding in the feature filter, and then calculate their cosine similarity. This measure reflects the learned semantic relationships between class labels. Figure 12 compares these three similarity measurements. All results indicate that Hydra automatically discovers both semantic and visual similarities between classes, with clearer distinctions between the super synsets compared to within them. This suggests that Hydra effectively captures the hierarchical structure of the classes, maintaining strong similarities within super synsets while preserving distinctions between them.
Figure 12: Cosine similarity between classes for gates, global features, and label vectors. Red numbers denote the average distances within each super synset region (snake, wading bird, furniture). All these measurements show that Hydra promotes similarity between classes within the same super synset while maintaining distinctions between different super synsets.

### v2

@@ Abstract
We present SimLAP, a universal contrastive learning framework for visual representations that extends conventional contrastive learning to accommodate arbitrary pairs. Our approach is validated using IN1K, where 1K diverse classes compose 500,500 pairs, most of them being distinct. Surprisingly, SimLAP achieves superior performance in this challenging setting. Additional benefits include the prevention of dimensional collapse and the discovery of class relationships. Our work highlights the value of learning common features of arbitrary pairs and potentially broadens the applicability of contrastive learning techniques on the sample pairs with weak relationships.
@@ 1 Introduction
Figure 1: Similarity distribution of class pairs under the subspaces of feature extracted by SimCLR and SimLAP for snake-lamp. We can hardly find any common visual features from the example of a garter snake and a table lamp, However, we find that the 500 dimensions with the lowest variance from SimCLR’s representation can separate snake-lamp from other classes. SimLAP learns such a subspace while representation learning.
@@
We present SimLAP by extending CL methods to accommodate arbitrary pairs, e.g., SimCLR. To show the robustness of dealing with disparate pairs, we train SimLAP on ImageNet (Deng et al., 2009), with 1,000 diverse classes forming 500,500 class pairs, most of which are disparate. The positive pairs are randomly selected from the 500,500 class pairs. Our model achieves superior performance in the challenging situation, demonstrating its robustness and effectiveness in discovering common features. In summary, our contributions are as follows:
  • We propose SimLAP a universal contrastive learning framework to allow learning visual representations from samples belonging to arbitrary class pairs through feature selection.
@@ 3.3 Universal Contrastive Learning
Figure 2: Universal contrastive learning for arbitrary class pairs. The feature filter generates a gate vector to activate the common features for the given class pair by averaging their label embeddings. SimLAP learns visual representations by maximizing the agreement of common features between disparate samples (Hydra-Lamp) in the corresponding subspace.
@@
The feature filter is a vital component of SimLAP to allow universal contrast learning. It controls the activation of each feature for a class pair y1-y2. The feature filter consists of two parts: label embedding and MLP layer. Label embedding converts discrete labels to continuous vectors. We use the mean of the two label vectors to represent the common information among a class pair. An MLP layer generates gate values to select the corresponding dimensions for the common information. The calculation is
@@ 4.1 Experimental Setting
All experiments are conducted using ResNet50 (He et al., 2016) as the backbone architecture, with ImageNet-1K (IN1K) as the training dataset. For baseline comparisons, we use pretrained models from official repositories: MoCov3, DINO, BarlowTwins, BYOL, and GPaCo; SimCLR implementation from VISSL (Goyal et al., 2021); Supcon trained for 600 epochs. Our proposed method is implemented in two variants: SimLAP is built upon SimCLR framework, Hydra-MoCo is built upon SimCLR framework MoCov3. Detailed training configurations and hyperparameters are provided in appendix D.
@@ 4.2 Main Result
We adopt the transfer learning performance with linear prob to evaluate the quality of learned representations. Table 1 compares UCL to ICL and CCL in terms of their transfer learning performance on six classification tasks: Cars, CIFAR10, CIFAR100, FLW, Pets, and STL, with the average (AVG) score also provided. For a fair comparison, we obtain two groups of results by using momentum or not, and pick their representatives: SimCLR as ICL and Supcon as CCL without momentum, MoCov3 as ICL and GPaCo as CCL with momentum. SimLAP demonstrate competitive performance in each group. Especially, SimLAP MoCo as well as MoCoV3 stands out with the highest average scores of 94.0%, indicating their superior performance across all tasks. Notablely, SimLAP MoCo is trained for 300 epochs, using only 30% of training for MoCov3. Overall, the table underscores the importance of uCL in achieving optimal transfer learning performance.
@@
Vision transformers (Dosovitskiy, 2020) are dominant in many CV tasks, especially beneficial from scaling in larger data (Dehghani et al., 2023). We study the effectiveness of our model for increasing the model size in table 2. We train 100 epochs for ViT-tiny, 300 epochs for ViT-small and ViT-base. SimLAP outperforms both supervised and unsupervised learning methods for the ViT-tiny model. The experimental results demonstrate that SimLAP scales well with increasing model size, outperforming or matching supervised learning across different ViT architectures. This scaling behavior, combined with Hydra’s ability to learn from arbitrary pairs, makes it a promising approach for leveraging large-scale, diverse datasets in visual representation learning tasks.
@@ 4.3 Universal Contrast Avoids Dimensional Collapse
To thoroughly investigate the dimensional collapse phenomenon, we construct IN1P, a focused subset of IN1K containing 12K images from 10 highly-related dog breeds. This smaller dataset allows us to extensively train on all possible class pairs, which is challenging in IN1K. We train models for 20,000 epochs to ensure comprehensive learning of all pairs. As shown in fig. 3(a), we compare the KNN classification performance (top-1 accuracy) on CIFAR10 across different methods. SimCLR exhibits clear signs of dimensional collapse after 3,000 epochs, evidenced by declining performance and decreasing singular values (Jing et al., 2022). This reveals that while the non-linear projector can delay dimensional collapse, it cannot prevent it entirely. In contrast, SimLAP demonstrates sustained performance improvement throughout the extended training period, surpassing both Supcon and SimCLR. This robustness against dimensional collapse can be attributed to our framework’s ability to leverage rich common features between classes.
@@
Figure 3: SimLAP prevents dimensional collapse and benefits from longer training. Trained on IN1P and evaluated on CIFAR10. Lower singular values suggest that the learned representations are concentrating information in fewer dimensions, that is, dimensional collapse.
@@ 4.4 Visualization
We analyze the quality of learned representations by examining embeddings from the backbone network (excluding the projector and filter). First, we compare the inter- and intra-class similarity distributions on the IN1K validation set. As shown in fig. 4, SimLAP achieves the sharpest inter-class similarity distribution and the smallest overlap between inter- and intra-class distributions compared to Supcon and SimCLR. This indicates that SimLAP learns representations that better preserve class-specific information while maintaining clear class boundaries.
To visualize these embedding properties, we apply $t$-SNE (Van der Maaten and Hinton, 2008) to the features of 10 classes from the IN1K validation set (fig. 5). The visualization reveals two key findings: (1) In the global space, SimLAP maintains clear class separation as good as the methods optimized in the global space, even through our model is optimized in subspaces; (2) When examining the subspace for specific class pairs (e.g., Garter snake-Chihuahua), our model successfully identifies shared features that bring these seemingly disparate classes closer while preserving the overall structure of the embedding space. This demonstrates our model’s unique ability to learn both class-discriminative features and shared characteristics across arbitrary class pairs.
Figure 4: Similarity distribution for SimLAP, Supcon, and SimCLR. The number in the middle denotes the overlap of two distributions. Smaller value means better class-separation.
@@
Figure 5: $t$-SNE visualization of learned embeddings for two contrasting groups: 5 dog breeds and 5 snake species. Large markers indicate class centers, with stars representing dog classes and points representing snake classes. While classes remain well-separated in the global space (abc), SimLAP can selectively bring disparate classes closer in their designated subspace (d) through learned feature filtering.
@@ 5.1 Ablation
Figure 8: Impact of semantic distance between positive pairs (N-closest classes, N=0 means same class). We report KNN classification accuracy on CIFAR10. While Supcon’s performance drops significantly with increasing semantic distance, SimLAP maintains robust performance through adaptively learning in subspaces.
@@
To systematically evaluate the limitations of supervised contrastive learning (Supcon) and identify scenarios where our method excels, we conduct experiments on IN25, a dataset of 25 classes with well-defined semantic relationships (details in section D.1). We systematically vary the semantic distance of positive pairs by composing them from N-closest classes, where N=0 represents pairs from the same class. As shown in fig. 8, we compare Supcon and SimLAP on CIFAR10 transfer learning performance as we increase the semantic distance between positive pairs. Supcon’s performance degrades consistently as the semantic distance increases, highlighting its limitation in learning from disparate pairs. In contrast, SimLAP maintains robust performance across different semantic distances, thanks to its ability to create appropriate subspaces for each pair.
@@ 5.2 Hyperparameter
We investigate the impact of each component in SimLAP. To evaluate the quality of learned representations, we apply the widely used linear prob (Chen et al., 2020) on IN1K, KNN (K=10) on CIFAR10 as a reference of transfer learning. SimLAP was trained on IN1K for 100 epochs. Table 3 demonstrates the effects of various hyperparameters. Our default settings for training are highlighted in Gray.
@@
We find that normalization at the end of the projector is critical to train SimLAP. Without normalization, the training process becomes unstable, resulting in NaN errors. Conversely, the model barely converges with Batch Normalization (Ioffe and Szegedy, 2015). This is because Batch Normalization enforces each dimension to be informative (high variance), conflicting with our filter module, which selectively closes certain dimensions. Our model only functions effectively with Layer Normalization  (Ba et al., 2016).
@@
Similar to other contrastive learning methods (Chen et al., 2020; He et al., 2020), SimLAP’s performance is influenced by the number of dimensions and the temperature parameter.
@@ 6 Conclusion
In this paper, we have reconsidered the concept of positive pairs in contrastive learning and explored the learning framework for arbitrary pairs. Our research breaks the limitation of positive pairs and demonstrates the potential of learning common features from seemly unrelated class pairs. The search space of positive pairs for different contrastive learning approaches. As the allowed distance for positive pairs increases, the search space expands exponentially. This expansion highlights the potential for more comprehensive and nuanced representation learning. Our approach, SimLAP, demonstrates remarkable robustness in dealing with arbitrary pairs, even in a challenging case where positive pairs are randomly picked and most of them are disparate. The method successfully learns transferable representations under these challenging conditions, validating our core idea.
@@ Appendix A Understanding from A Case Study
Considering the large number and diversity of classes in IN1K, we use a case study to demonstrate how SimLAP interprets the similarity between classes. Specifically, we select 17 classes belonging to the super synset ‘snake.n.01’, 16 classes belonging to the super synset ‘wading bird.n.01’, and 21 classes belonging to the super synset ‘furniture.n.01’ from the IN1K validation set. Each class contains 50 images. This case study demonstrates that SimLAP promotes
We summarise the gate values of SimLAP in Figure 9. Specifically, we pass 1K classes to the filter to generate 1000 vectors with 256-D gate values, which indicate the activation of the corresponding dimensions for each class. Each light point in the heatmap represents the activation status of a specific dimension for the corresponding class. From the figure, we can observe that the gates exhibit a roughly binary behavior (either activated or not), and the activation vectors for different classes are distinct. This indicates that each class creates a unique subspace within the feature space. These experimental observations support our hypothesis that SimLAP can effectively learn to create class-specific subspaces.
@@
We investigate how SimLAP identifies common features between similar classes within a coarse category. For the subclasses within a coarse class, the common features are dominant, i.e. showing high similarity in the global space, as discussed earlier. To delve deeper, we select the top-10 dimensions with the highest activation for subclasses within each coarse class, as illustrated in Figure 11. These dimensions likely represent the most salient features for distinguishing the subclasses. We then calculate the class similarity of features in the subspace defined by these dimensions for each coarse class. Figure 10 visualizes the class similarity in the subspaces computed for three coarse classes: Snake, Bird, and Furniture. Compared to the class similarity in the global space, the discriminative information is strengthened for the corresponding subspace. Specifically, the gap between intra and inter coarse-class similarity is magnified to 0.5 for Snake in the Snake-Snake Subspace, 0.7 for Bird in the Bird-Bird Subspace, and 0.5 for Furniture in the Furniture-Furniture Subspace. The features in the subspaces must contain distinct information about the class pair. These features also contain visual information about other classes such that these classes are still discriminative in the unrelated subspaces. These results demonstrate SimLAP’s ability to create meaningful subspaces that amplify the differences between the classes, even when no common features appear to exist in the global space, like Furniture-Snake.
Figure 10: Class similarity in subspaces. The gap between intra and inter coarse-class similarity is widened in their corresponding subspaces, highlighting SimLAP’s effectiveness in creating discriminative subspaces.
@@
We investigate three similarity measurements to evaluate the ability to identify the relationships between classes: 1) Dot product of gates: We calculate the gates for each class and sharpen the values to 1 if above 0.5 and to 0 otherwise. This measure indicates how similar the activated dimensions are between classes, reflecting the structural similarity of their subspaces. 2) Features in the global space: For each class pair, we use the global features to calculate the cosine similarity between all sample pairs coming from each class pair. We then use the average to denote the similarity between classes. This represents the overall visual similarity in the original feature space. 3) Class similarity in label embedding: We obtain the class vectors through the label embedding in the feature filter, and then calculate their cosine similarity. This measure reflects the learned semantic relationships between class labels. Figure 12 compares these three similarity measurements. All results indicate that SimLAP automatically discovers both semantic and visual similarities between classes, with clearer distinctions between the super synsets compared to within them. This suggests that SimLAP effectively captures the hierarchical structure of the classes, maintaining strong similarities within super synsets while preserving distinctions between them.
Figure 12: Cosine similarity between classes for gates, global features, and label vectors. Red numbers denote the average distances within each super synset region (snake, wading bird, furniture). All these measurements show that SimLAP promotes similarity between classes within the same super synset while maintaining distinctions between different super synsets.